# Eigenspectrum Analysis of Neural Networks without Aspect Ratio Bias

### v1

@@ Abstract
Diagnosing deep neural networks (DNNs) through the eigenspectrum of weight matrices has been an active area of research in recent years. At a high level, eigenspectrum analysis of DNNs involves measuring the heavytailness of the empirical spectral densities (ESD) of weight matrices. It provides insight into how well a model is trained and can guide decisions on assigning better layer-wise training hyperparameters. In this paper, we address a challenge associated with such eigenspectrum methods: the impact of the aspect ratio of weight matrices on estimated heavytailness metrics. We demonstrate that matrices of varying sizes (and aspect ratios) introduce a non-negligible bias in estimating heavytailness metrics, leading to inaccurate model diagnosis and layer-wise hyperparameter assignment. To overcome this challenge, we propose FARMS (Fixed-Aspect-Ratio Matrix Subsampling), a method that normalizes the weight matrices by subsampling submatrices with a fixed aspect ratio. Instead of measuring the heavytailness of the original ESD, we measure the average ESD of these subsampled submatrices. We show that measuring the heavytailness of these submatrices with the fixed aspect ratio can effectively mitigate the aspect ratio bias. We validate our approach across various optimization techniques and application domains that involve eigenspectrum analysis of weights, including image classification in computer vision (CV) models, scientific machine learning (SciML) model training, and large language model (LLM) pruning. Our results show that despite its simplicity, FARMS uniformly improves the accuracy of eigenspectrum analysis while enabling more effective layer-wise hyperparameter assignment in these application domains. In one of the LLM pruning experiments, FARMS reduces the perplexity of the LLaMA-7B model by 17.3% when compared with the state-of-the-art method.
@@ 1 Introduction
Figure 1: Comparing FARMS with common HT-SR methods in analyzing the weight matrices of a DNN. We use the weight matrix $\mathbf{W}_{i}$ of a well-trained layer with a large aspect ratio as an example (i.e., $\gamma=m/n$ is much larger than 1). Due to the influence of the aspect ratio $\gamma=m/n\gg 1$, the ESD is more concentrated and less HTed than other layers. As a result, previous methods for fitting a power-law distribution (shown as “PL Fitting”) tend to overestimate the heavytailness metric PL_Alpha_Hill and incorrectly suggest that the layer is poorly trained. Due to this aspect ratio bias, some optimization methods (such as TempBalance) cannot accurately assign per-layer hyperparameters, leading to suboptimal training. In contrast, our method, FARMS, which conducts ESD analysis with a fixed aspect ratio, accurately measures the training quality of the layer. See a concrete numerical example in Figure 7 in Appendix A.
@@
Random Matrix Theory (RMT) has long been used as a foundational tool in multiple research domains (Guhr et al., 1998; Tulino and Verdú, 2004; Bai and Silverstein, 2010; Tao, 2012; Nadakuditi and Newman, 2012) and has now been applied to providing precise characterizations of neural networks (NNs) (Pennington and Worah, 2017; Dobriban and Wager, 2018; Adlam and Pennington, 2020; Mei and Montanari, 2022; Hastie et al., 2022; Hu and Lu, 2022; Ba et al., 2022; Couillet and Liao, 2022; Derezinski et al., 2020). Among several emerging topics in this field, Heavy-Tailed Self-Regularization (HT-SR) (Martin and Mahoney, 2021; Mahoney and Martin, 2019; Martin and Mahoney, 2019; Martin et al., 2021) Theory has been gaining significant attention. Unlike conventional RMT approaches, HT-SR Theory focuses on studying weight matrices of DNNs with strongly correlated elements, a typical characteristic of well-trained, practical DNNs. The essence of HT-SR theory is that the weight matrices of well-trained DNNs are strongly correlated, and that leads to heavy-tailed (HTed) ESDs (Martin and Mahoney, 2021). Analyzing these HT ESDs provides valuable insights into the quality of trained models, which provides methods and techniques for model diagnostics and model training. For example, Martin and Mahoney (2021) show that one can locate undertrained layers by finding weight matrices with less HTed ESDs, which provides useful metrics for model ranking, comparison and selection (Martin et al., 2021; Yang et al., 2023). Furthermore, one can use a larger learning rate on the undertrained layers (Zhou et al., 2024), which provides an empirically successful training method to balance these layers with other, more well-trained ones. This technique has been developed into a general approach to improve test accuracy, training efficiency, and model interpretability in various applications, such as image classification (Zhou et al., 2024), LLM model pruning (Lu et al., 2024), and SciML model training and fine-tuning (Liu et al., 2024, 2025b).
@@
The core of the HT-SR theory lies in the measurement of the heavytailness of ESDs. In this context, we identified an often-overlooked issue in prior studies: weight matrices of different sizes are analyzed using the same HT measuring procedure, even though their spectral distributions can differ in theory. Specifically, algorithms inspired by the HT-SR theory often begin by measuring the ESD of a weight matrix $\mathbf{W}$, that is, the empirical distribution of eigenvalues of the correlation matrix $\mathbf{W}^{\top}\mathbf{W}$. The ESD will then be analyzed and used for downstream applications such as model selection, training, and compression (Mahoney and Martin, 2019; Martin and Mahoney, 2019, 2021; Zhou et al., 2024; Lu et al., 2024; Liu et al., 2024). For weight matrices initialized using the i.i.d. random Gaussian elements, it is well-known that the ESD converges to the Marchenko–Pastur (MP) distribution as the dimensions of the matrix grow infinitely large in the proportional limit, i.e., when the number of rows $m$ and the number of columns $n$ increase but the ratio $m/n$ stays fixed. The MP distribution depends explicitly on the aspect ratio $\gamma=m/n$, since this aspect ratio determines the limiting bulk range, $[(1-\sqrt{\gamma})^{2},(1+\sqrt{\gamma})^{2}]$, and it influences the overall shape of the ESD. As a result, weight matrices of different aspect ratios exhibit distinct shapes in their ESDs, especially when they transition from the i.i.d. initialization to being correlated. Comparing ESDs with different aspect ratios $\gamma$ directly, as is commonly done in prior work, overlooks this dependency on aspect ratio and can lead to inaccurate conclusions. Therefore, a careful consideration of the aspect ratio is critical when analyzing or comparing weight matrices, especially in eigenspectrum analysis.
@@
Considering these challenges, we introduce FARMS (Fixed-Aspect-Ratio Matrix Subsampling), a new method for measuring HT characteristics in HT-SR weight analysis. FARMS analyzes the training quality of layers by partitioning each weight matrix into (overlapping) submatrices of fixed aspect ratios and then do an eigenspectrum analysis on the average ESD of these submatrices. This approach enables accurate computation of HT-SR metrics regardless of variations in matrix aspect ratios, ensuring a robust evaluation of layer quality across diverse matrix sizes that one can have in a DNN.
@@
To validate the effectiveness of FARMS, we conduct experiments across various application domains, including CV, SciML, and LLM pruning. Additionally, these experiments are conducted across different parameter settings and model architectures. We compare FARMS with several prior HT-SR approaches that use weight eigenspectrum analysis for hyperparameter scheduling (Zhou et al., 2024; Lu et al., 2024; Liu et al., 2024). We also do weight analysis to measure various post-training and pruned models, making FARMS useful for efficient, compressed models. Our findings demonstrate that models optimized using FARMS exhibit lower mean and variation in HT-SR metrics across layers, a sign of good-quality training as reported in prior work (Martin et al., 2021; Liu et al., 2024). Our code is available here [^1] https://github.com/HUST-AI-HYZ/FARMS. Our key contributions are summarized below.
@@
  • Mitigating Aspect Ratio Bias in Eigenspectrum Analysis. FARMS addresses the aspect ratio bias of existing HT-SR eigenspectrum analysis, and it enables better computation of HT metrics on various DNN models. This improvement is achieved through a subsampling-based HT estimation method independent of the aspect ratio of weight matrices. In particular, we use a numerical example in Figure 6 and multiple real-data experiments in Section 4.3 to demonstrate that FARMS mitigates the aspect ratio bias in training.
@@
  • Improved Layer-wise Hyperparameter Tuning. Since HT-SR Theory has been recently applied to various layer-wise hyperparameter tuning methods, FARMS thus improves these methods by offering a more accurate evaluation of HT metrics. In particular, we conduct experiments on DNN training and pruning across various model architectures. In CV models like ResNet and VGG, integrating FARMS into learning rate assignments yields improved accuracy compared to TempBalance, a recently proposed layer-wise learning rate scheduling method (Zhou et al., 2024). In LLMs pruning, FARMS improves AlphaPruning (Lu et al., 2024), the SOTA method in assigning layer-wise pruning ratios. Specifically, FARMS can reduce the perplexity of the LLaMA-13B model from 2029.20 to 413.76 using the magnitude pruning method at a 0.7 sparsity ratio. As another example in LLaMA-7B, it reduces the perplexity from 96.02 to 79.42 using the SparseGPT pruning method (Frantar and Alistarh, 2023) at a 0.8 sparsity ratio. In SciML, FARMS helps scientific models achieve up to a 5.66% error reduction during fine-tuning compared to the HT-SR based method TB_Sigmoid (Zhou et al., 2024), even at relatively low L2 relative error (L2RE) levels.
@@ 2.1 Prior Work on HT-SR Theory
In this section, we provide an overview of prior work on HT-SR theory, a framework derived from RMT and statistical physics that is relevant to understanding modern, practical DNNs. HT-SR theory (Martin and Mahoney, 2021, 2019) originates from RMT but extends well beyond it. The theory was proposed based on the observation that well-trained, state-of-the-art DNNs often exhibit HTed structures in the ESD of each layer. In the meantime, several rigorous theories in SGD relating HT phenomena to generalization performance were established, providing further theoretical support for HT-SR theory (Hodgkinson and Mahoney, 2021; Hodgkinson et al., 2022; Gurbuzbalaban et al., 2021; Simsekli et al., 2019, 2020). Building on the theoretical foundation of HT-SR, a model analysis tool called WeightWatcher (Martin and Mahoney, 2021) was developed. Without accessing any training or test data, methods based on HT-SR Theory can be used to assess the performance of models across various domains, such as CV and NLP (Martin and Mahoney, 2021; Martin et al., 2021; Yang et al., 2023; Liu et al., 2025b).
@@ 2.2 Optimization Methods Based on Eigenspectrum Analysis
Our paper is mainly motivated by HT-SR theory, focusing on the eigenspectrum analysis of weight matrices. Our paper connects to several DNN optimization methods that use eigenspectrum analysis to improve model performance. For example, spectral norm regularization has been applied to improve generalizability of trained models (Yoshida and Miyato, 2017; Miyato et al., 2018; Farnia et al., 2018). Stable rank normalization (SRN) (Sanyal et al., 2019), a normalization technique, scales matrices using their stable rank to enhance training stability in GANs and generalization in DNNs. TE-NAS (Chen et al., 2021) is a training-free neural architecture search framework that identifies high-performing networks by analyzing the neural tangent kernel spectrum and input space linear regions.
@@ 3.1 Typical HT-SR Analysis
Before introducing FARMS, we first review the commonly adopted procedures in HT-SR weight analysis. HT-SR analysis relies on estimating the layer quality based on the HT characteristic of the layer ESDs, which is quantified by HT metric PL_Alpha. The ESD of a weight matrix is the histogram of eigenvalues. The ESD of weight matrices evolves during training, transitioning from a bulk-dominated regime to an HTed regime Martin et al. (2021). The HTed portion can be modeled by a power-law (PL) distribution within an interval ($\lambda_{\text{min}}$, $\lambda_{\text{max}}$):
@@
where $\alpha$, the PL exponent, is a critical metric for analyzing training quality. To fit a PL distribution to the ESD, methods in HT-SR often use the Hill Estimator (Hill, 1975; Zhou et al., 2024; Lu et al., 2024; Liu et al., 2024)  [^2] One important point to note is that estimating PLs is inherently a challenging task (Clauset et al., 2009). Previous research suggests using the maximum likelihood estimate (Alstott et al., 2014; Martin and Mahoney, 2021). However, empirical evidence indicates that the Hill estimator performs more reliably in DNN optimization applications (Zhou et al., 2024; Lu et al., 2024; Liu et al., 2024).. For the $i$-th layer, suppose the weight matrix is $\mathbf{W}_{i}$ and the correlation matrix $\mathbf{W}_{i}^{\top}\mathbf{W}_{i}$ has ascending eigenvalues $\{\lambda_{i}\}_{i=1}^{n}$. The Hill estimator calculates PL_Alpha_Hill as:
[…]
where $k$ is an adjustable parameter. Changing $k$ essentially changes the lower eigenvalue threshold $\lambda_{\text{min}}$ for (truncated) PL estimation. Various metrics have been proposed to analyze the properties of ESDs, among which shape metrics—those that characterize the distributional shape of ESDs—have been shown to effectively predict the training quality of individual layers. The PL_Alpha_Hill metric is one such shape metrics.
There are several metrics used to quantify the structure of ESDs in HT-SR theory. In this work, we mainly consider the PL_Alpha_Hill, which is empirically shown to be effective for training tasks (Zhou et al., 2024; Lu et al., 2024; Liu et al., 2024; Qing et al., 2024). In general, undertrained layers in DNNs tend to exhibit larger PL_Alpha_Hill values, whereas well-trained or over-trained layers typically have smaller PL_Alpha_Hill values. This metric can also be used for a comprehensive analysis across a series of DNN models with similar architectures to find the best model. Well-trained models tend to exhibit a lower average PL_Alpha_Hill across all layers. Moreover, in fine-tuning tasks, a larger amount of data generally results in a lower standard deviation (STD) of PL_Alpha_Hill across the model, indicating a more balanced training progression across different layers (Liu et al., 2024).

### v2

@@ Abstract
Diagnosing deep neural networks (DNNs) by analyzing the eigenspectrum of their weights has been an active area of research in recent years. One of the main approaches involves measuring the heavytailness of the empirical spectral densities (ESDs) of weight matrices. This analysis has been shown to provide insights to help diagnose whether a model is well-trained or undertrained, and has been used to guide training methods involving layer-wise hyperparameter assignment. In this paper, we address an often-overlooked challenge in estimating the heavytailness of these ESDs: the impact of the aspect ratio of weight matrices. We demonstrate that matrices of varying sizes (and aspect ratios) introduce a non-negligible bias in estimating the heavytailness of ESDs, leading to inaccurate model diagnosis and layer-wise hyperparameter assignment. To overcome this challenge, we propose FARMS (Fixed-Aspect-Ratio Matrix Subsampling), a method that normalizes the weight matrices by subsampling submatrices with a fixed aspect ratio. Instead of measuring the heavytailness of the original ESD, we measure the average ESD of these subsampled submatrices. We show that this method effectively mitigates the aspect ratio bias. We validate our approach across various optimization techniques and application domains that involve eigenspectrum analysis of weights, including image classification in computer vision (CV) models, scientific machine learning (SciML) model training, and large language model (LLM) pruning. Our results show that despite its simplicity, FARMS uniformly improves the accuracy of eigenspectrum analysis while enabling more effective layer-wise hyperparameter assignment. In one of the LLM pruning experiments, FARMS reduces the perplexity of the LLaMA-7B model by 17.3% when compared with state-of-the-art methods.
@@ 1 Introduction
Figure 1: Comparing FARMS with common HT-SR methods in analyzing the weight matrices of a DNN. We use the weight matrix $\mathbf{W}_{i}$ of a well-trained layer with a large aspect ratio as an example (i.e., $\gamma=m/n$ is much larger than 1). Due to the influence of the aspect ratio $\gamma=m/n\gg 1$, the ESD is more concentrated and less HTed than other layers. As a result, previous methods for fitting a power-law distribution (shown as “PL Fitting”) tend to overestimate the heavytailness metric PL_Alpha (the HT index) and incorrectly suggest that the layer is poorly trained. Due to this aspect ratio bias, some optimization methods (such as TempBalance) cannot accurately assign per-layer hyperparameters, leading to suboptimal training. In contrast, our method, FARMS, which conducts ESD analysis with a fixed aspect ratio, accurately measures the training quality of the layer. See a concrete numerical example in Figure 7 in Appendix A.
@@
Random Matrix Theory (RMT) has long been used as a foundational tool in multiple research domains (Guhr et al., 1998; Tulino and Verdú, 2004; Bai and Silverstein, 2010; Tao, 2012; Nadakuditi and Newman, 2012) and has now been applied to providing precise characterizations of neural networks (NNs) (Pennington and Worah, 2017; Dobriban and Wager, 2018; Adlam and Pennington, 2020; Mei and Montanari, 2022; Hastie et al., 2022; Hu and Lu, 2022; Ba et al., 2022; Couillet and Liao, 2022; Derezinski et al., 2020). Among several emerging topics in this field, Heavy-Tailed Self-Regularization (HT-SR) (Martin and Mahoney, 2021; Mahoney and Martin, 2019; Martin et al., 2021) Theory has been gaining significant attention. Unlike conventional RMT studies, HT-SR Theory focuses on studying weight matrices with strongly correlated elements, a typical characteristic of well-trained, practical DNNs. These strong correlations lead to heavy-tailed (HTed) ESDs (Martin and Mahoney, 2021). Analyzing these HT ESDs provides valuable insights into the quality of trained models, enabling both model diagnostics and improvements in model training. For example, Martin and Mahoney (2021) show that one can locate undertrained layers by finding weight matrices with less HTed ESDs, which provides useful metrics for model selection and ranking (Martin et al., 2021; Yang et al., 2023). Furthermore, one can use a larger learning rate on the undertrained layers (Zhou et al., 2024), which provides an empirically successful training method to balance these layers with other, more well-trained ones. This technique has been developed into a general approach to improve test accuracy, training efficiency, and model interpretability in various applications, such as image classification (Zhou et al., 2024), LLM model pruning (Lu et al., 2024), and SciML model training and fine-tuning (Liu et al., 2024, 2025b).
@@
The core of the HT-SR theory lies in the measurement of the heavytailness of ESDs. In this context, we identified an often-overlooked issue in prior studies: weight matrices of different sizes are analyzed using the same HT measuring procedure, even though their ESDs can differ in theory. Specifically, algorithms inspired by the HT-SR theory often begin by measuring the ESD of a weight matrix $\mathbf{W}$, that is, the empirical distribution of eigenvalues of the correlation matrix $\mathbf{W}^{\top}\mathbf{W}$. The shape of the ESD is then analyzed by fitting it to certain distribution families, such as the power-law (PL) distribution. For weight matrices initialized using the i.i.d. random Gaussian elements, it is well-known that the ESD converges to the Marchenko–Pastur (MP) distribution as the dimensions of the matrix grow infinitely large in the proportional limit, i.e., when the number of rows $m$ and the number of columns $n$ increase but the ratio $m/n$ stays fixed. The MP distribution depends explicitly on the aspect ratio $\gamma=m/n$, since this ratio determines the limiting bulk range, $(1-\sqrt{\gamma})^{2},(1+\sqrt{\gamma})^{2}]$, and it influences the overall shape of the ESD. As a result, weight matrices of different aspect ratios exhibit distinct shapes in their ESDs, especially when they transition from the i.i.d. initialization to being correlated. Comparing ESDs with different aspect ratios $\gamma$ directly, as is commonly done in prior work, overlooks this dependency on aspect ratio and can lead to inaccurate shape estimates. Therefore, a careful consideration of the aspect ratio is critical when analyzing or comparing weight matrices, especially in eigenspectrum analysis.
@@
Considering these challenges, we introduce FARMS (Fixed-Aspect-Ratio Matrix Subsampling), a new method for measuring HT characteristics. FARMS partitions each weight matrix into (overlapping) submatrices of fixed aspect ratios and then does an eigenspectrum analysis on the average ESD of these submatrices. This approach enables accurate computation of heavytailness metrics regardless of variations in matrix aspect ratios, ensuring a robust evaluation of layer quality across diverse matrix sizes that one can have in a DNN.
@@
To validate the effectiveness of FARMS, we conduct experiments across various application domains, including CV, SciML, and LLM pruning. In addition, these experiments are conducted on different parameter settings and model architectures. We compare FARMS with several prior HT-SR approaches that use weight eigenspectrum analysis for layer-wise hyperparameter assignment (Zhou et al., 2024; Lu et al., 2024; Liu et al., 2024). We also apply FARMS to measure various post-training and pruned models, making FARMS useful for model compression. Our findings demonstrate that models optimized using FARMS exhibit lower mean and variation in HT-SR metrics across layers, a sign of good-quality training as reported in prior work (Martin et al., 2021; Liu et al., 2024). Our code is available here [^1] https://github.com/HUST-AI-HYZ/FARMS. Our key contributions are summarized below.
@@
  • Mitigating Aspect Ratio Bias in Eigenspectrum Analysis. FARMS addresses the aspect ratio bias of existing HT-SR eigenspectrum analysis, and it enables better computation of HT metrics on DNN models that have heterogeneous layer types. This improvement is achieved through a subsampling-based HT estimation method independent of the aspect ratio of weight matrices. In particular, we use a numerical example in Figure 6 and multiple real-data experiments in Section 4.3 to convincingly demonstrate that FARMS mitigates the aspect ratio bias in training.
@@
  • Improved Layer-wise Hyperparameter Tuning. Since HT-SR Theory has been recently applied to various layer-wise hyperparameter tuning methods, FARMS thus improves these methods by offering a more accurate evaluation of HT metrics. In particular, we conduct experiments on DNN training and pruning across various model architectures. In CV models like ResNet and VGG, integrating FARMS into learning rate assignments yields improved accuracy compared to TempBalance, a recently proposed layer-wise learning rate assignment method (Zhou et al., 2024). In LLMs pruning, FARMS improves AlphaPruning (Lu et al., 2024), the SOTA method in assigning layer-wise pruning ratios. Specifically, FARMS can reduce the perplexity of the LLaMA-13B model from 2029.20 to 413.76 using the magnitude pruning method at a 0.7 sparsity ratio. As another example in LLaMA-7B, it reduces the perplexity from 96.02 to 79.42 using the SparseGPT pruning method (Frantar and Alistarh, 2023) at a 0.8 sparsity ratio. In SciML, FARMS helps scientific models achieve up to a 5.66% error reduction during fine-tuning compared to the HT-SR based method TB_Sigmoid (Zhou et al., 2024), even at relatively low L2 relative error (L2RE) levels.
@@ 2.1 Prior Work on HT-SR Theory
In this section, we provide an overview of prior work on HT-SR theory, a framework derived from RMT that is relevant to understanding modern, practical DNNs. HT-SR theory (Martin and Mahoney, 2021; Mahoney and Martin, 2019) originates from RMT but extends well beyond it. The theory was proposed based on the observation that well-trained, state-of-the-art DNNs often exhibit HTed structures in the ESD of each layer. In the meantime, several rigorous theories in SGD relating HT phenomena to generalization performance were established, providing further theoretical support for HT-SR theory (Hodgkinson and Mahoney, 2021; Hodgkinson et al., 2022; Gurbuzbalaban et al., 2021; Simsekli et al., 2019, 2020). Building on the theoretical foundation of HT-SR, a model analysis tool called WeightWatcher (Martin and Mahoney, 2021) was developed. Without accessing any training or test data, methods based on HT-SR Theory can be used to assess the training quality of models across various domains, such as CV and NLP (Martin and Mahoney, 2021; Martin et al., 2021; Yang et al., 2023; Liu et al., 2025b).
@@ 2.2 Optimization Methods Based on Eigenspectrum Analysis
Our paper is mainly motivated by HT-SR theory, focusing on the eigenspectrum analysis of weight matrices. Our paper connects to several DNN optimization methods that use eigenspectrum analysis to improve model performance. For example, spectral norm regularization has been applied to improve generalizability of trained models (Yoshida and Miyato, 2017; Miyato et al., 2018; Farnia et al., 2019). Stable rank normalization (SRN) (Sanyal et al., 2020), a normalization technique, scales matrices using their stable rank to enhance training stability in GANs and generalization in DNNs. TE-NAS (Chen et al., 2021) is a training-free neural architecture search framework that identifies high-performing networks by analyzing the neural tangent kernel spectrum and input space linear regions.
@@ 3.1 Typical HT-SR Analysis
Before introducing FARMS, we first review the commonly adopted procedures in HT-SR weight analysis. HT-SR analysis relies on estimating the layer quality based on the HT characteristic of the layer ESDs, which is quantified by HT metric PL_Alpha, introduced below. The ESD of a weight matrix is the histogram of eigenvalues. The ESD of weight matrices evolves during training, transitioning from a bulk-dominated regime to an HTed regime Martin et al. (2021). The HTed portion can be modeled by a power-law (PL) distribution within an interval ($\lambda_{\text{min}}$, $\lambda_{\text{max}}$):
@@
where $\alpha$, the PL exponent, is a critical metric for analyzing training quality. To fit a PL distribution to the ESD, methods in HT-SR often use the Hill Estimator (Hill, 1975; Zhou et al., 2024; Lu et al., 2024; Liu et al., 2024)  [^2] One important point to note is that estimating PLs is inherently a challenging task (Clauset et al., 2009). Previous research suggests using the maximum likelihood estimate (Alstott et al., 2014; Martin and Mahoney, 2021). However, empirical evidence indicates that the Hill estimator performs more reliably in DNN optimization applications (Zhou et al., 2024; Lu et al., 2024; Liu et al., 2024).. For the $i$-th layer, suppose the weight matrix is $\mathbf{W}_{i}$ and the correlation matrix $\mathbf{W}_{i}^{\top}\mathbf{W}_{i}$ has ascending eigenvalues $\{\lambda_{i}\}_{i=1}^{n}$. The Hill estimator calculates PL_Alpha_Hill (the PL exponent PL_Alpha estimated using the Hill estimator) as:
$\texttt{PL\_Alpha\_Hill}=1+\frac{k}{\sum_{i=1}^{k}\ln\frac{\lambda_{n-i+1}}{\lambda_{n-k}}},$ (2)
where $k$ is an adjustable parameter. Changing $k$ essentially changes the lower eigenvalue threshold $\lambda_{\text{min}}$ for (truncated) PL estimation. Various metrics have been proposed to analyze the properties of ESDs, among which shape metrics, which characterize the distributional shape of ESDs, have been shown to effectively predict the training quality of individual layers (Yang et al., 2023). The PL_Alpha_Hill metric is one such shape metrics.
There are several metrics used to quantify the structure of ESDs in HT-SR theory. In this work, we mainly consider the PL_Alpha_Hill, which is empirically shown to be effective for training tasks (Zhou et al., 2024; Lu et al., 2024; Liu et al., 2024; Qing et al., 2024). In general, undertrained layers in DNNs tend to exhibit larger PL_Alpha_Hill values, whereas well-trained or over-trained layers typically have smaller PL_Alpha_Hill values. This metric can also be used for a comprehensive analysis across a series of DNN models with similar architectures to find the best model. Well-trained models tend to exhibit a lower average PL_Alpha_Hill across all layers (Martin et al., 2021; Yang et al., 2023). Moreover, in fine-tuning tasks, a larger amount of data generally results in a lower standard deviation (STD) of PL_Alpha_Hill across the model, indicating a more balanced training progression across different layers (Liu et al., 2024).